Write the rate expression (differential rate 

equation) for the following reaction:

CH4(g) + 2O2(g) =2(g) + 2H2O(g)


1
Expert's answer
2021-01-11T03:47:38-0500

rate = - d[CH4]/dt = -1/2 x d[O2]/dt = +1/2 x d[CO2]/dt = +1/2 x d[H2O]/dt